Solve:
6^log base 6 (10^x)=100
Help please
thank you

Answer by dabanfield(803)

6^log base 6 (10^x)=100
Since by definition log (base 6) a = b if and only if 6^b = a the left side becomes 10^x so we have:
10^x = 100
Taking the base 10 log of each side above we have:
log (base 10) 10^x = log (base 10) 100
x = 2